OpenStudy (anonymous):

4. What is the slope of the line represented by the equation 10x + 5y = 4 ?

OpenStudy (anonymous):

a) 2 b) 1/2 c) -1/2 d) -2

OpenStudy (studygurl14):

10x + 5y = 4 to put it in y = mx + b 5y = -10x + 4 y = -2x + 4/5 m = slope thus, -2 = m